answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of flipping a coin 4 times and getting heads each time?

The probability on the first flip is 50% .

The probability on the 2nd flip is 50% .

The probability on the 3rd flip is 50% .

The probability on the 4th flip is 50% .

The probability of 4 heads is (50% x 50% x 50% x 50%) = (0.5)4 = 1/16 = 6.25%

What step in the risk management process is focused on determining the probability and severity of a hazard?

Assess hazards is the step in the composite risk management process that is focused on determining the probability and severity of a hazard occurring.

A can be any What a chance a person takes without knowing what the outcome will be the degree of probability or likelihood that something will happen a dangerous chanc?

Anyone who is going to answer this will need to know what choices there are for the answers. Without knowing this they will not be able to know which cause something dangerous to happen.

Comparision between binomial plus normal plus hypergeometric distribution?

The binomial distribution is a discrete probability distribution which describes the number of successes in a sequence of draws from a finite population, with replacement. The hypergeometric distribution is similar except that it deals with draws without replacement. For sufficiently large populations the Normal distribution is a good approximation for both.

If 63 out of 287 select a and 23 out of 301 select b then how many times is a more likely to be selected?

a is selected by 63 out of 287 - a rate of 63/287.

b is selected at a rate of 23/301.

Selection rate for a/selection rate for b = (63/287)/(23/301) = 0.2195/0.0764 = 2.8727

Ie a is likely to be selected 2.8727 times as often as b, or a is 1.8727 times more likely to be selected.

If you toss a penny how do you think the coin will land heads or tails explain?

there is even chance of having heads or tails since there is only 1 head and 1 tail on a penny.

What is the probability of flipping NO heads when you flip 3 identical coins?

The chance of not flipping a head in each instance is 1/2. You need that to happen three times.

1/2 x 1/2 x 1/2 = 1/8

So there is a 1 in 8 chance of getting no heads from 3 coin flips.

What is the probability of rolling a sum which is a multiple of five?

There are 36 permutations of rolling two standard dice. Of them, four (1+4, 2+3, 3+2, and 4+1) sum to five, and three (6+4, 5+5, and 4+6) sum to ten. The probability, then of rolling a multiple of five is 7 in 36, or about 0.1944.

How many ways are there to arrange the letters in the word ADDRESS if the two Ds must be together?

The permutations of the letters ADDRESS if the two D's must be together is the same as the permutations of the letters ADRESS, which is 6 factorial, or 720, divided by 2, to compensate for the two S's, which means that the number of distinct permutations of the letters ADDRESS, where the two DD's must be together is 360.

How many card that is not a diamond face card?

There are 52 cards in a deck. 52cards/4 suits = 13. 13x3=39. There are 39 cards in a deck that are not diamond.

When apply geomatric distribution?

When you want to find the probability of something happening for the first time while repeating an experiment, you use the geometric distribution. For example, you are throwing a six-faced die and you are expecting the number '6' to be thrown for the first time.

A definition and an example of multiple representation?

it's basically higher order thinking ex) 3+4-2*4dived by 8= e

Trending Questions
From a deck of 52 playing cards what is the probability that the card is a queen or a red card? In an experiment h is the control variable k is the response variable and l and m are extraneous factors. The specific units of are referred to as the treatments of the experiment? Are you likely to be weary after a rogaining event? What is the single parameter in the binomial distribution and what sample statistic would be used to estimate it? Is drawing marbles from a bag without replacing them dependent or independent? What is the probability of getting either a sum of 5 or at least one 5 in the roll of a pair of dice? How do you arrange the word kctrice? How are books classified? Probability of rolling a 1 and a 1? What is the probability of getting an even number if the die is rolled once? Can you have a period 4 days after conception? What is it called when you get a probability that is 0? How many ways can Cindy arrrange 42 books in different shelf if each shelf has to be the same? What does it mean when the standard error value is smaller than the standard deviation? In a third grade class containing 12 girls and 10 boys Two students are selected randomly to give out papers What is the probability that the second student chosen is a boy if the first was a girl? If a coin is flipped 7 times what is the outcomes? What is the difference between sampling frame and sampling uinit? When the population standard deviation is unknown the sampling distribution is equal to what? There are 12 students in a classroom including the triplets Joey Chloe and Zoe. If 3 of the 12 are randomly selected to give speeches what is the probability that Joey will be first Chloe will be seco? How many ways is there to seat six guest at a table for 6?